ICM 1998, Berlin, Aug. 18–27

Abstracts of Plenary and Invited Lectures

Section:

7. Lie Groups and Lie Algebras

1991 MS Classification: 22 Kottwitz, Robert, University of Chicago, USA Harmonic analysis on semisimple p-adic Lie algebras

Certain topics in harmonic analysis on semisimple groups arise naturally when one uses the Arthur-Selberg trace formula to study automorphic representations of adele groups. These topics, which fall under the heading of “comparison of orbital integrals, ” have been surveyed in Waldspurger’s article in the proceedings of ICM94. As in Harish-Chandra’s work, many questions in harmonic analysis on the group can be reduced to analogous questions on its Lie algebra, and in particular this is the case for “comparison of orbital integrals. ” We will discuss some recent work of this type.

ICM 1998 Berlin · Plenary and Invited Lectures